OpenStudy (anonymous):

A flying bird drops a seed. The heights, h, in meters, of the seed above the ground can bemodelled by the relation h=-5r^2 + 125 there t is in seconds. How far above the ground is the bird when it drops the seed? and how long does the seed take to hit the ground?

OpenStudy (anonymous):

is that \(\large h=-5\color {red}t^2+125 \) ??

OpenStudy (anonymous):

set \[-5r^2+125=0\] and solve for \(r\)

OpenStudy (anonymous):

oh sorry yes. it's not r its t

OpenStudy (anonymous):

h= -5t^2 +125

OpenStudy (anonymous):

or set \(\large-5\color {red}t^2+125=0\) and solve for \(\color{blue}{t}\)

OpenStudy (anonymous):

\[5\color{pink}t^2=125\] \[\color{purple}{t}=25\] \[\color{green}{t}=5\]

OpenStudy (anonymous):

okay so thats the seconds. how would I see how far above the ground is the bird when it dropped the seed?

OpenStudy (anonymous):

WAIT, I sub in 25 for t and calculate it right?

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{n}}\color{orange}{\normalsize\text{o}}\color{#9c9a2e}{\normalsize\text{t}}\color{green}{\normalsize\text{ }}\color{blue}{\normalsize\text{t}}\color{purple}{\normalsize\text{o}}\color{purple}{\normalsize\text{o}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{h}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{r}}\color{blue}{\normalsize\text{d}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{r}}\color{red}{\normalsize\text{i}}\color{orange}{\normalsize\text{g}}\color{#9c9a2e}{\normalsize\text{h}}\color{green}{\normalsize\text{t}}\color{blue}{\normalsize\text{?}}\color{purple}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

im kinda lost. how do I calculate the height of the bird when the seed was dropped? can you show me the steps?

OpenStudy (anonymous):

no the bird drops the seed at time \(t=0\) and when \(t=0\) you get \(h=125\)

OpenStudy (anonymous):

soooo .......how did you get that t=0

OpenStudy (anonymous):

well the problem is not really presented that well is it? it just gives you an equation and does not say why

OpenStudy (anonymous):

i guess i have to say i know it because i know it, which admittedly is not a good answer but \(t\) represents time in seconds, and you have to start counting some time. so it should really say "a bird drops a seed and after that the height is given by \(h(t)=-5t^2+125\) so written in to the problem is that you start counting when the bird drops the seed, that is when \(t=0\)

OpenStudy (anonymous):

sorry i could not come up with a better answer, it is a rather artificial question and in fact the equation is not even right. they should have simply asked "what is \(h(t)=-5t^2+125\) when \(t=0\) ?"and "solve \(-5t+125=0\)" and let you get on with it

OpenStudy (anonymous):

okay thanks :)